Amarillo Slim, a professional dart player, has an 80% chance of hitting the bull's eye on a dartboard with any throw. Assume each throw at the dartboard is independent of any other throw.

Suppose, now that Amarillo Slim throws 10 darts, one at a time, at the dartboard.  What is the probability he gets exactly 7 bull's eyes? (to 4 decimals)
Answer #1

Solution

Given that ,

p = 0.80

q = 1 - p = 1 - 0.80 = 0.20

n = 10

Using binomial probability formula ,

P(X = x) = (n C x) * p x * (1 - p)n - x

P(X = 7) = (10 C 7) * 0.80 7 * (0.20)10 - 7

Probability =0.2013